Run Planning Meeting Notes, Nov 8-Nov 14, 2018
From GlueXWiki
<- Back to Run Coordination Meetings:Fall2018 Run
- Trigger graphs
- Trigger Spreadsheet
- Low energy running Trigger Spreadsheet
- Run Plan Google Calendar
- Accel. Plan Google Calendar
- Acc whiteboard
- Counting house whiteboard
- Hall D schedule for week 8
- Hall D schedule for week 9
Contents
[hide]Wednesday, Nov 14 at 08:45
- Last 24h:
- Installation of beam pipe
- Harp scan [1] [2] [3] [4]: Proposal for convergence
- Switch back to standard production running:
- Ramp up PS magnet to 910A (done)
- Low-energy tagger counters off (done)
- Set coherent peak at 8.8GeV (done)
- Issues
- Reports
- Monitoring (Thomas)
- Rates (Alexandre)
- ComCal (Sasha)
- Plans
- Beam tuning
- Production running
- TAC run: Thursday while Sasha is on shift?
- Notes
Tuesday, Nov 13 at 08:45
- Last 24h:
- 6h+ scheduled down time due to system recovery, but many problems
- Preparations for beam pipe installation (done)
- Upgrade to Hall D current locks + beam based test (30min)
- Only very little beam during the night, accelerator harp scan done [5][6], ours interrupted
- Switch back to standard production running:
- Ramp up PS magnet to 910A (done)
- Low-energy tagger counters off (done)
- Set coherent peak at 8.8GeV (not done)
- Issues
- Reports
- Monitoring (Thomas)
- Rho polarization for low-energy run (Alex)
- Plans
- Beam Studies (Tuesday 08:30, 8h): installation of beam pipe
- French delegation to visit ComCal (6 people)
- Harp scan
- Production running
- Notes
Monday, Nov 12 at 08:45
- Last 72h:
- 6h down time on Friday due to HV issue on RF cavity 1L23
- 12h down on Saturday due to 1L23 and more
- 4h down on Sunday
- Completed the goal of 5 full cycles (2B events) on Sunday ~5pm
- Issues
- Reports
- Plans
- System Recovery (Monday 07:30, 6h+): switching back to 8.8GeV running?
- Beam Studies (Tuesday 08:30, 8h): installation of beam pipe
- Notes
Friday, Nov 9 at 08:45
- Last 24h:
- >80% beam available, >95% used
- completed 2nd cycle
- harp scan ongoing
- Issues
- Sluggish CSS during the day, investigating
- Tagger dump BPM AD00c
- Reports
- Plans
- Beam Studies and System Recovery switched
- System Recovery (Monday 08:30): investigate water leak in tunnel, ~3h
- Beam Studies (Tuesday): high current test, halls A, B and C have to stayed locked, D optional
- Notes
Thursday, Nov 8 at 08:45
- Last 24h:
- TAC run with empty target, v-wire at 50nA, reduced field in PS (400A), low E TAGH counters on
- Harp Scan
- Beam tuning (deferred)
- Data taking with reduced PS field, coherent edge at 7GeV
- TAGH Rates (Amo), (Diamond), Coincidence Rates
- Low energy running Trigger Spreadsheet
- RF problems in the night
- Issues
- PS and PSC coincidence scalers frozen, ps_client running on gluon100 restarted, instructions in hdops wiki
- beam position went awry, rates soared, reference module for IOC reading out 5C11B failed, replaced
- EPICS froze for 2h, started working without intervention, now again unresponsive
- Target PTR turned off, do not ignore any alarms!
- Reports
- Plans
- Notes